OpenStudy (anonymous):

A skateboarder, starting from rest, rolls down a 13.5-m ramp. When she arrives at the bottom of the ramp her speed is 7.45 m/s. (a) Determine the magnitude of her acceleration, assumed to be constant. m/s2 (b) If the ramp is inclined at 22.0° with respect to the ground, what is the component of her acceleration that is parallel to the ground? m/s2

OpenStudy (amistre64):

might want to us some energy formulas

OpenStudy (amistre64):

potential energy at the top of the hill is equal to the kinetic energy at the bottom P = K mgh = 1/2 mv^2 or gh = v^2/2 but this is related to newtons second law by F = ma, so gh = a